Use the Substitution method to solve the system of equations.
x + y = 10
y = x + 8

Answer by checkley77(12844)

x + y = 10
x=10-y
y = x + 8
y=(10-y)+8
y=10-y+8
2y=18
y=18/2
y=9 answer.
x+9=10
x=10-9
x=1 answer.
proof:
9=1+8
9=9